OpenStudy (kmullis6):

Has anyone taken algebra 1 on flvs?

OpenStudy (anonymous):

FLVS?

OpenStudy (jskhupmang):

yes

OpenStudy (kmullis6):

florida virtual school @Chiko_1278

OpenStudy (animelover98):

Yes

OpenStudy (anonymous):

no but i go on K12

OpenStudy (kmullis6):

have you done 2.04 module 2 quiz? @AnimeLover98 @jskhupmang

OpenStudy (kmullis6):

I'm very confused about everything, and need help solving the problems.

OpenStudy (anonymous):

yea

OpenStudy (animelover98):

Yes, I can help you, but I refuse to give answers.

OpenStudy (anonymous):

no i have connections academy

OpenStudy (kmullis6):

@AnimeLover98 No, I don't want the answers, just help.

OpenStudy (animelover98):

What is your question?

OpenStudy (kmullis6):

@Spamchopper7 So the assignments are the same?

OpenStudy (yanasidlinskiy):

@kmullis6 a more proper place for this kind of question is in the chats. As we all know, post your question and let someone answer it. If you don't get it, then you can tag someone.

OpenStudy (kmullis6):

Angelica's service club is selling flowers for a fundraiser. Each flower costs them $0.50 to purchase. They set the sale price of each flower at the purchase price plus 40%. Write and solve an equation to determine how many flowers they will need to sell to make a profit of $500. Use complete sentences to explain what each number and variable in your equation represents. Write your solution in the context of the original problem.

OpenStudy (animelover98):

P = the profit they make n = the number of flowers they sell m = the mark-up of the flowers c = the cost of a flower

OpenStudy (animelover98):

Can you fill in the variables for this equation? n = P / (mc - c)

OpenStudy (kmullis6):

Sadly, no. I'm horrible at algebra, and it being online doesn't make it any easier to retain. /:

OpenStudy (animelover98):

The P or Profit would be the $500 they need

OpenStudy (animelover98):

Each flower is $.50, which would be the cost. 40% is the mark-up. The number they will sell is n.

OpenStudy (animelover98):

So, what does the equation look like?

OpenStudy (kmullis6):

Would it be like n = 500 for the first part? I'm probably wrong lol

OpenStudy (animelover98):

P=500 n is unknown. We are trying to solve for n. n is the amount of flowers.

OpenStudy (animelover98):

Here, I'll start you out. n=500/(mc - c)

OpenStudy (animelover98):

Find what numbers go in (mc - c)

OpenStudy (kmullis6):

Um.. 40%-0.50 ?

OpenStudy (animelover98):

YES! Exactly! (1.4*0.5 - 0.5) Make sure you change 40% to 1.4

OpenStudy (kmullis6):

Oh ok!

OpenStudy (animelover98):

So what is your equation?

OpenStudy (animelover98):

It should look like this n = 500 / (1.4*0.5 - 0.5)

OpenStudy (kmullis6):

Ok. We still need to find n right?

OpenStudy (animelover98):

After you solve this equation, the answer will be n. Solve for the parentheses first.

OpenStudy (kmullis6):

I got 1/5

OpenStudy (animelover98):

You should get n=500/5 Solve that.

OpenStudy (kmullis6):

100

OpenStudy (animelover98):

Whoops, sorry I typed the wrong thing in. You should have n = 500 / (1.4*0.5 - 0.5) n = 2500

OpenStudy (kmullis6):

Ohh ok.

OpenStudy (animelover98):

So, when you get 1/5, that is equal to .20

OpenStudy (animelover98):

Just divide .20 by 500, and you get 2500.

OpenStudy (animelover98):

The number of flowers they need to sell is 2500 to make a profit of $500. That is, if they sell their flowers at $.50 marked up to 40%.

OpenStudy (animelover98):

Good job working through this!

OpenStudy (kmullis6):

Sorry I didn't respond eariler, I had to step away from the computer. Thank you so much, I really appreciate it! (:

OpenStudy (animelover98):

Of course! Happy to help. I totally understand how you feel, I struggle with math really badly. In fact, I'm doing my Algebra 2 Honors course on FLVS right now haha.